OpenStudy (anonymous):

A square with a side of 6 is decreased to a side of 4. By how much is the diagonal decreased? square with side 4 aligned with square of side 6 at lower left vertex 2 times the square root of 2 3 times the square root of 2 2 times the square root of 3 3 times the square root of 3 A square with a side of 6 is decreased to a side of 4. By how much is the diagonal decreased? square with side 4 aligned with square of side 6 at lower left vertex 2 times the square root of 2 3 times the square root of 2 2 times the square root of 3 3 times the square root of 3 @Mathematics

OpenStudy (anonymous):

the original diagonal was \[6\sqrt{2}\] and the new one is \[4\sqrt{2}\]

OpenStudy (anonymous):

subtract to get \[2\sqrt{2}\]

OpenStudy (anonymous):

thank u :)

OpenStudy (anonymous):

yw